OpenStudy (anonymous):

If the radius of the Earth is roughly 3,960 miles, how many times larger is the volume of the Earth than the volume of a ping-pong ball? A ping-pong ball has a radius of 0.7441 inches. 1 mile = 5,280 feet.

OpenStudy (aum):

Volume of a sphere = \(\large \frac 43 \pi r^3\) \[ \frac{\text{Volume of Earth}}{\text{Volume of ping-pong ball}} =\large \frac{\frac 43 \pi r_e^3}{\frac 43 \pi r_p^3} = \frac{r_e^3}{r_p^3} \\ r_e = \text{Radius of the Earth} = 3960 \times 5280 \times 12 ~\text{inches.} \\ r_p = \text{Radius of the ping-pong ball} = 0.7441 ~\text{inch.} \]Plug in the numbers and calculate the ratio.